OpenStudy (anonymous):

Word Problem: A ball is thrown upwards. The height h, measured in feet, attained by it after being in the air for t seconds follows the rule h= -16t^2+32t+50. From what height was the ball thrown upward?

OpenStudy (anonymous):

set t to 0 and you'll have the answer

OpenStudy (anonymous):

since at the moment it is thrown it is not yet in the air, and since time started counting as soon as the ball is thrown in the air, t=0 is the exact moment the ball is thrown.

OpenStudy (anonymous):

and thus, the height that will be the result of that equation is the height at which it is thrown.

OpenStudy (anonymous):

Lol. Lets work this out together. I understand this more with actual numbers.

OpenStudy (anonymous):

ok

OpenStudy (snowfire):

zimmah is right, this equation relates the time and height of the ball, so when you just throw the ball, it will be t=0

OpenStudy (snowfire):

So plug in 0 for every t you see, and there you go.

OpenStudy (anonymous):

\[h=-16*0^2+32*0+50 = ?\]

OpenStudy (anonymous):

Yeah, Im understanding him. I just want to clarify it more. So Im replacing all of t with 0. Im going to get 50 as an answer.

OpenStudy (anonymous):

@zimmah h=50?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it's good to ask when you need more clarification, the purpose of asking is to understand it completely. you can never ask too many questions.

OpenStudy (anonymous):

Sweet. So if the ball was at 30feet above the ground ... how would I set that up for a solution?

OpenStudy (anonymous):

You know how to find the time at which the ball crosses the x-axis right?

OpenStudy (anonymous):

Like for example if i would say \[ax^2+bx+c=0\] you would be able to solve that, right?

OpenStudy (anonymous):

Yeah, isn't that the form for Quadratic Equation?

OpenStudy (anonymous):

correct. so let's use that knowledge, with a light alteration, to find what we need to know

OpenStudy (anonymous):

because we do not want to know when it hits the ground, we want to know when it hits a 'imaginary ground' so to speak, that is 30 feet above it, and runs paralel to it.

OpenStudy (anonymous):

so we are going to just say the ground is 30 feet above itself by saying \[ax^2+bx+c=30\] or in your case \[30= -16t^2+32t+50\]

OpenStudy (anonymous):

however there is a catch

OpenStudy (anonymous):

the quadratic equation only works when the other side is 0. so we have to make it 0 again by moving the 30 to the other side first.

OpenStudy (anonymous):

so it becomes -30 and goes to the other side so that we get \[-16t^2+32t+50-30=0\]

OpenStudy (anonymous):

Awwwh !!! So therefore 30 - 50 .

OpenStudy (anonymous):

Wait. Wait. Let me solve. on paper. Ill tell you the answer I get.

OpenStudy (anonymous):

ok, however make sure you use 50-30. (I may have confused you a bit by having the equation the other way around)

OpenStudy (anonymous):

50-30 will give 20. So now I have: 0 = -16t^2 + 32t + 20

OpenStudy (anonymous):

Im going to do the Quadratic Equation now.

OpenStudy (anonymous):

yes, that's right.

OpenStudy (anonymous):

So i solved and got: \[\frac{ -32+/- [48] }{ -32 } \]

OpenStudy (anonymous):

oh hey

OpenStudy (anonymous):

lets see

OpenStudy (anonymous):

ye, we can even simplyfy it a bit more to two readable answers.

OpenStudy (anonymous):

you know, the +/- means you get two answers, one for the+and one for the -

OpenStudy (anonymous):

note that this not necessarily means you get a positive and a negative end results, it's just the number under the root that can be both positive and negative

OpenStudy (anonymous):

\[\frac{ -32+48 }{ -32 }\] and \[\frac{ -32-48 }{ -32 }\]

OpenStudy (anonymous):

I got these: \[\frac{ 16 }{ -32 } and \frac{ -80 }{ -32 } = \frac{ 80 }{ 32 }\]

OpenStudy (anonymous):

yes, these numbers are correct still, yes you can still simplify them.

OpenStudy (anonymous):

you can easily tell if you can simplify a number by breaking it into prime numbers. Prime numbers are the 'atoms' of the numbers so to speak. It's pretty interesting because every number has a unique signature of prime numbers.

OpenStudy (anonymous):

usually it's enough to know the first few, 2,3,5,7,9,11,13 you just try to divide until you can no longer divide and eventually you will end up with the DNA of the number

OpenStudy (anonymous):

now why would you want to do that, you may ask?

OpenStudy (anonymous):

\[\frac{ 16 }{ -32 }\frac{ 2*2*2*2 }{ -2*2*2*2*2 }=\frac{ 2^4 }{ -2^5 }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

I simplified all the way and got this: \[-\frac{ 16 }{ 32 } =\frac{ 1 }{ 2 }\] and \[\frac{ 80 }{ 32 } = \frac{ 5 }{ 2 }\]

OpenStudy (anonymous):

It felt like I did something wrong. =\

OpenStudy (anonymous):

no you did right, you have the right solution, except you forgot the - sign at the 1/2

OpenStudy (anonymous):

why does it feel wrong?

OpenStudy (anonymous):

Because the number jumped from 80/32 to 5/2 , haha. Just I thought the number was too low, a bit of curiosity.

OpenStudy (anonymous):

lol i forgot the - sign at the last one as well, we both did, that's quite funny.

OpenStudy (anonymous):

oh well let me show you

OpenStudy (anonymous):

look \[80=40*2=20*2^2=10*2^3=5*2^4\]

OpenStudy (anonymous):

\[32=16*2=8*2^2=4* 2^3=2^5\]

OpenStudy (anonymous):

\[\frac{ 2^4 }{ 2^4 }=\frac{ 1 }{ 1 }=1\]

OpenStudy (anonymous):

anything times 1 is still the same, so we can drop the 1

OpenStudy (anonymous):

Interesting!! , I was not looking at it that way. But thats explainable! The answer to the question would be: At points 1/2, 5/2, the ball will be 30 feet above the ground.

OpenStudy (anonymous):

that's what i said, you can do this for ANY number that is not a prime number.

OpenStudy (anonymous):

you basically take the number, and split it into the prime numbers it is made of, i always call it the number DNA.

OpenStudy (anonymous):

you can most easily do it by starting with the lowest prime numbers but in some cases another prime number or even a multiple of several prime numbers may be easier.

OpenStudy (anonymous):

Sweet! Hey would you be down for just one more problem?

OpenStudy (anonymous):

for example if a number ends with a 0 you can always divided by 2 and 5 (often multiple times)

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

A ball was thrown vertically upwards from top of a 25 foot tower with an initial velocity of 45 feet per second. How high did the ball go?

OpenStudy (anonymous):

ah, that's a physics formula

OpenStudy (anonymous):

No wonder I purposely skipped it. I did everything else but that.

OpenStudy (anonymous):

what's the gravity constant in feet? i only know the one for meters.

OpenStudy (anonymous):

I have no clue. I am not in Psychics.

OpenStudy (anonymous):

i will google

OpenStudy (anonymous):

0.304800 let's just make it 0.3

OpenStudy (anonymous):

doesn't matter that much for the purpose of explaining anyway

OpenStudy (anonymous):

so, we have a few clues in the text.

OpenStudy (anonymous):

first, the height at time 0 is 25.

OpenStudy (anonymous):

Yepp and the the initial velocity is 45 feet per second.

OpenStudy (anonymous):

beat me to it

OpenStudy (anonymous):

now, it's not explicitly given but, we assume there is gravity there.

OpenStudy (anonymous):

and since we defined gravity as 0.3 ft/s^2 (it's an acceleration, not a velocity)

OpenStudy (anonymous):

and since the are also no other forces active (for the sake of simplicity we will ignore drag because drag will make your head spin)

OpenStudy (anonymous):

we can make a formula for the height using these pieces of knowledge.

OpenStudy (anonymous):

Thats a lot of new information. =\ , But go along. Im totally following you.

OpenStudy (anonymous):

the +25 is pretty easy to explain, because the 25 at the start is not a function of time (otherwise it would be 0 and cancel out at time 0)

OpenStudy (anonymous):

so, we write h=25 but obviously we are still missing some parts otherwise the ball would freeze in mid air an forever stay at 25 ft

OpenStudy (anonymous):

so we need a function of time as well.

OpenStudy (anonymous):

which is the speed initially given to it (45ft/s) so, every second of t, the ball moves 45ft

OpenStudy (anonymous):

so after t, it moved 45, after 2t it moved 90 and so on

OpenStudy (anonymous):

so what we have so far is \[h(t)=45t+25 (still~ incomplete!) \]

OpenStudy (anonymous):

you understand it so far?

OpenStudy (anonymous):

Lol. Im following let me soak it all in. Im re-reading everything. Standby.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Alrighty. So how would t = 25?

OpenStudy (anonymous):

in the function we have till now? it's not complete yet you know that right?

OpenStudy (anonymous):

\[h(25)=45*(25)+25\]

OpenStudy (anonymous):

which is 1150 btw

OpenStudy (anonymous):

H = 46 ?

OpenStudy (anonymous):

What do you mean by that, you want to know when H reaches 46?

OpenStudy (anonymous):

Oh because h(25) = 25h So in the equation its like: 25h = 1150 divide both sides by 25.

OpenStudy (anonymous):

no you misinterpret that

OpenStudy (anonymous):

Officially, when written like h(t)=something the h(t) means that h is a function of t

OpenStudy (anonymous):

in this case, height is a function of time. (we want to know which height the ball has at any given time, that is the purpose of the function, to draw a graph of h against t)

OpenStudy (anonymous):

so the h(25) in this case does not mean 25h but it simply means the value of t that h is a function of is 25.

OpenStudy (anonymous):

in other words it sais, for every single t in the equation, replace it with 25. so you can ignore the numbers on the right.

OpenStudy (anonymous):

Oh okay! I get it. Its just like f(x)